OpenStudy (anonymous):

simplify z^2-4/2z^2+5z+2 x 2z^2+z/z-2

OpenStudy (phi):

\[ \frac{z^2-4}{2z^2+5z+2} \cdot \frac{2z^2+z}{z-2}\] You should factor the polynomials. a difference of squares (a^2 - b^2) factors into (a-b)(a+b) in 2z^2 + z you can factor out a z